# A Laboratory Method for Measuring the Cross-Polarization in High-Contrast Imaging

### v1

@@ Abstract
Electric Field Conjugation (EFC) and related techniques have proven to be effective for high-contrast coronagraphic imaging, and they have applications in current and future exoplanet direct imaging missions. EFC utilizes one or more deformable mirrors (DMs) to nullify the electric fields from the host star within a designated “dark hole” region in the image plane, allowing for the detection of much fainter planetary emissions. The quality of a dark hole is quantified by the contrast, i.e., the planet-to-star intensity ratio, that could be measured inside it. Inside the dark hole, the planet’s light is distinguished from the residual starlight due to its incoherence; specifically, unlike starlight, the planetary light is essentially unmodulated by DM perturbations. Cross polarization can cause the modulation of the starlight to behave differently than predicted, and can limit the achievable contrast ratio, especially if it is not taken into account. The cause of cross polarization is the various weakly polarizing interactions of the light as it propagates through the optical system, notably reflections and finite focal length effects. Sophisticated physical optics software can model the cross polarization for a for a fully specified optical system, but this capability does not obviate the need for end-to-end measurements of the cross polarization.
This article presents realistic numerical simulations of a novel laboratory-based method for measurements of the cross-polarization in a coronagraphic optical system. This capability may well prove useful for validating models of polarization effects in future high-contrast imaging systems. The proposed laboratory method is not much more demanding than current efforts on the high-contrast testbeds, yet the simulation results demonstrate highly accurate estimates of the real and imaginary parts of the electric field corresponding to the cross polarization in a dark hole. These encouraging results suggest that laboratory implementation should be viable.
The proposed method uses a 100 mW supercontinuum laser with a narrow bandpass and two linear polarizers to isolate the cross polarization effects. A novel nonlinear probing scheme, along with nonlinear regression, is introduced to address polarizer leakage. Simulations are performed for an aberrated Lyot coronagraph with an f-number of 44, at a wavelength $\lambda=1\,\mu$m, using a single $33\times 3$ 33 actuator deformable mirror. The initial dark hole contrast is $\sim 10^{-10}$, with cross-polarization electric fields corresponding to intensities of $\sim 10^{-11}$ in contrast units. These cross-polarization electric fields are estimated simultaneously across 441 pixels, covering a patch approximately $4\lambda/D\times 4\lambda/D$ in the image plane.
The primary application of this method is likely laboratory validation of digital twin models for cross-polarization. On-sky, the experimentally validated models could then be used to account for cross polarization when estimating the planetary signal. Although speculative, an exciting future prospect is experimental validation of AI-based models that leverage conventional EFC sensing data to predict cross-polarization electric fields.
@@ 1 Introduction
This article presents a concept for a relatively simple laboratory experiment and DM-based sensing algorithms that will allow DM modulation-based measurements of the electric fields corresponding to the polarization effects. The most likely use of these laboratory measurements is the validation of models of the cross-polarization, which is timely given the increasing importance of digital twin concepts in high-contrast imaging.[10] At the simplest level, the proposed measurements could validate basic descriptive statistics (e.g., mean and variance) of the cross-polarization. More sophisticated would be validating AI-based models that map measurements of the fields made in course of classical EFC procedures to the cross-polarized electric fields.
1.1 Examples of Cross-Polarization in Simple Focusing Systems
@@ 1.1.1 Simulation Examples
In the case of Fig. 1, “ideal” only means that the focusing is aberration-free (despite the extreme $f$/#). Although such a small $f$/# would be useful in confocal microscopy, it is not useful for telescopes; however, it provides a helpful and relatively simple illustration of the cross-polarization phenomenon. Fig. 2 shows the corresponding magnitudes $|\mathrm{fields}|$ in the focal plane, with the top row corresponding to $x$-polarization in the input beam, and $y$-polarization in the bottom row. It is important to notice that $|E_{x}|$ in the top row is the same as $|E_{y}|$ in the bottom row; these are the dominant polarizations defined in Sec. 1.2. Similarly, $|E_{y}|$ in the top row is the same as $|E_{x}|$ in the bottom row; these are the cross-polarizations also defined in Sec. 1.2. The analytical developments to follow will take advantage of this symmetry.
@@
1.2 Calculation of Intensities Including Cross-Polarization
@@
2 Including Two Linear Polarizers
@@
  • The linear polarizer in the entrance beam only allows A-polarized light to pass (see Sec.1.2). This results in the dominant polarization being in the $x$-direction at the detector plane and the cross-polarization being in the $y-$direction [see Eq. (9)].
@@
3 Scalar Fourier Optics Approximations and Symmetry Considerations
@@
All of the calculations presented in this article are carried out without paraxial assumptions. The simulation results in Figs. 2 and 3 described above in Sec. 1.1.1 as well as those from the OAP-based coronagraph simulated in detail (described later) exhibited two symmetries for the dominant and cross fields. Interestingly, even beams off-axis by angles $100\lambda/D$ also showed these same symmetries. The following equations express the two symmetries:
@@ 5 The Dark Hole
6 Using Probes for Focal Plane Wavefront Estimation
@@ 6.1 Regression Equations
Measuring the intensity with $P$ different probe commands, $\boldsymbol{c}_{1},\,\dots,\,\boldsymbol{c}_{P}$ enables the regression on the real and imaginary parts of the cross field in the dark hole, $\boldsymbol{f}_{y}(\boldsymbol{0})$. The corresponding dominant field $\boldsymbol{f}_{x}(\boldsymbol{0})$ could be estimated as well (with a sufficient number of probes), but the linear polarizers make this difficult, so it is better to estimate it without the polarizers in place. In the Discussion section, we outline a practical approach to adjust the dark hole command to account for the wavefront error introduced by the polarizers. For simplicity, we consider only estimating the real and imaginary parts of the cross field, $\boldsymbol{f}_{y}(\boldsymbol{0})$. When we apply the linear polarizers as described in Sec. 2, the measured intensity in the detector pixels with probe command $\boldsymbol{c}_{i}$ is:
@@ 7 Simulations and Results
All optical propagation computations were carried out using the VirtualLab Fusion (VLF) software from LightTrans, Inc. The simulation results in this article correspond to a square Lyot-type stellar coronagraph shown schematically in Fig. 7 and with key parameters summarized in Table 1. This coronagraph consists of 3 OAPs. The initial propagation direction of the beam before it encounters OAP1 is in the $+z$ direction. The simulation is configured so that the principal axis of the beam remains in the $x-z$ plane after each OAP encounter. With this arrangement, the $y-$axis is invariant throughout the system. Assuming no polarizers in the system, the consequence of this invariance is that $y-$polarized light in the input beam results in $y-$polarized light in the detector plane, weak polarization effects due to the finite $f$/# not withstanding. Similarly, if the light in the input beam is $x-$polarized it must also be $x-$polarized in the detector plane since it must be orthogonal to the global $y-$ and local $z-$directions (again, weak polarization effects due to the finite $f$/# not withstanding). Thus, the $A$ and $B$ directions discussed in Sec. 1.2 correspond directly to the $x$ and $y$ axes in the input beam. To construct the $\tilde{\boldsymbol{D}_{x}}$ and $\tilde{\boldsymbol{D}_{y}}$ matrices introduced in Sec. 4, VLF carried out 1089 propagations-one for each DM actuator-through the optical system. For each of these propagations, the input beam was $x-$polarized and set to one of the DM influence functions with an amplitude of unity. In the detector plane, the corresponding dominant field values, corresponding to a column of $\tilde{\boldsymbol{D}_{x}}$, were the “$E_{x}$” values of the VLF output at each pixel, and the cross field values, corresponding to a column of $\tilde{\boldsymbol{D}_{y}}$, were the “$E_{y}$” values.

### v2

@@ Abstract
Electric Field Conjugation (EFC) and related techniques have proven to be effective for stellar coronagraphy, but existing methods are not suited for treating small polarization effects, referred to here as cross polarization, resulting from propagation through the optical system. EFC utilizes a deformable mirror (DM) to nullify the electric fields from the host star within a dark hole region in the image plane, allowing for the detection of much fainter planets.
This article presents realistic numerical simulations of a laboratory method for measuring the cross-polarization electric field in a stellar coronagraph. This novel method uses linear polarizers to isolate the cross polarization component and probing procedures similar to those in existing EFC methods to measure the electric field corresponding to cross polarization. The proposed nonlinear probing scheme is specifically designed to address polarizer leakage. The method is not a generalization of existing EFC procedures in order to null the cross polarization field; rather, it is a non-iterative laboratory calibration procedure with applications in instrument characterization and validation of numerical models. However, lab-validated numerical models of the cross-polarization are directly applicable to on-sky observations.
The proposed experimental method is not significantly more demanding than existing testbed procedures, requiring little additional equipment. Yet, the simulation results demonstrate highly accurate estimates of the electric field corresponding to the cross polarization in a dark hole. These encouraging results suggest that laboratory implementation is viable. Simulations are performed for a randomly aberrated Lyot coronagraph with a dark hole size of $\sim 4\lambda/D\times 4\lambda/D$ in the image plane with a contrast of $\sim 10^{-10}$ in contrast (i.e., the planet-to-star intensity ratio) units. The cross-polarization electric fields accurately estimated by the method correspond to intensities of $\sim 10^{-11}$ in contrast units.
@@ 1 Introduction
1.1 Summary of Article Content
This article presents simulations of a concept for a relatively simple, yet highly sensitive, laboratory method for measuring the electric fields corresponding to the contribution of the cross polarization to the total intensity in a dark hole. Its most likely uses are laboratory calibration of cross polarization and validation of cross polarization models, which is timely given the increasing importance of digital twin concepts in high contrast imaging.[10] At the simplest level, the proposed measurements could validate models of basic descriptive statistics (e.g., mean and variance) of the cross-polarization intensity. More sophisticated would be validating AI-based models that map measurements of the dominant field made in the course of classical EFC procedures to the cross-polarization field (“cross field” for short). Lab-validated models of the cross field would potentially improve the achievable contrast in on-sky observations because polarization effects could be taken into account in the estimate of the planetary signal or possibly even nulled, but these aspects are not the subject of this article.
While this method uses probing and concepts that are similar to those in the EFC sensing step, it is critical to understand that this method is not a generalization of EFC. This method only measures the cross field, and it is not iterative. Assuming access to a high-contrast testbed with a coronagraph and working EFC procedures to create a dark hole, the additional lab equipment required includes a 100 mW supercontinuum laser with a continuous wave (CW) mode and two linear polarizers with less than $10^{-5}$ leakage of the orthogonal intensity. The proposed procedure consists of the following steps, which are explained in more detail in the course of this article:
  1. 1.
    Without the linear polarizers in place, apply EFC procedures to make a dark hole. Take note of the DM command corresponding to the dark hole, $\boldsymbol{c}_{0}$.
  2. 2.
    Insert the two linear polarizers. One goes just after the entrance pupil and the other goes just before the detector. They are arranged so that if there were no polarizing effects in the optical system, all of the light would be extinguished, leakage from the polarizers notwithstanding. This arrangement isolates the cross polarization arriving at the detector. Note that isolating the cross polarization in this way comes at the expense of extinguishing any planetary light, which does not conflict with the objective of measuring the cross field. (Indeed, one is free to remove the polarizers after making the measurements.)
  3. 3.
    Apply the probing procedures detailed in this article to measure the intensities needed to feed the regression method.
  4. 4.
    Apply the regression method detailed in this article to estimate the electric field associated with the cross polarization in the dark hole.
Sec. 2 provides simulation examples of cross-polarization arising in simple focusing optical systems. Sec. 3 elucidates the relationship between coherence, polarization and intensity, as applied in this article. Sec. 4 presents the matrix-based coronagraph models underpinning the numerical methods presented here and introduces the dark hole. Sec. 5 details the probing and analysis procedures for estimating the cross fields while accounting for leakage of orthogonal intensity from the linear polarizers. Sec. 6 describes the numerical simulations of the proposed method and the resulting estimates of the cross field. Finally, Sec. 7 reviews key points of this work and discusses various practical aspects of carrying out the lab experiment, namely laser fluctuations, stray light from the polarizers and wavefront error introduced by the polarizers. In addition, it outlines how one might make use of a lab-validated model of the cross field on-sky.
2 Examples of Cross-Polarization in Simple Focusing Systems
@@ 2.1 Simulation Examples
In the case of Fig. 1, “ideal” only means that the focusing is aberration-free (despite the extreme $f$/#). Although such a small $f$/# would be useful in confocal microscopy, it is not useful for telescopes; however, it provides a helpful and relatively simple illustration of the cross-polarization phenomenon. Fig. 2 shows the corresponding magnitudes $|\mathrm{fields}|$ in the focal plane, with the top row corresponding to $x$-polarization in the input beam, and $y$-polarization in the bottom row. It is important to notice that $|E_{x}|$ in the top row is the same as $|E_{y}|$ in the bottom row; these are the dominant polarizations defined in Sec. 3. Similarly, $|E_{y}|$ in the top row is the same as $|E_{x}|$ in the bottom row; these are the cross-polarizations also defined in Sec. 3. The analytical developments to follow will take advantage of this symmetry.
@@
3 Calculation of Intensities Including Cross-Polarization
@@
3.1 Including Two Linear Polarizers
@@
  • The linear polarizer in the entrance beam only allows A-polarized light to pass (see Sec.3). This results in the dominant polarization being in the $x$-direction at the detector plane and the cross-polarization being in the $y-$direction [see Eq. (9)].
@@
3.2 Scalar Fourier Optics Approximations and Symmetry Considerations
@@
All of the calculations presented in this article are carried out without paraxial assumptions. The simulation results in Figs. 2 and 3 described above in Sec. 2.1 as well as those from the OAP-based coronagraph simulated in detail (described later) exhibited two symmetries for the dominant and cross fields. Interestingly, even beams off-axis by angles $100\lambda/D$ also showed these same symmetries. The following equations express the two symmetries:
@@ 4.1 The Dark Hole
5 Estimation of the Cross-Polarization Electric Field
@@ 5.1 Regression Equations
Measuring the intensity with $P$ different probe commands, $\boldsymbol{c}_{1},\,\dots,\,\boldsymbol{c}_{P}$ enables the regression on the real and imaginary parts of the cross field in the dark hole, $\boldsymbol{f}_{y}(\boldsymbol{0})$. The corresponding dominant field $\boldsymbol{f}_{x}(\boldsymbol{0})$ could be estimated as well (with a sufficient number of probes), but the linear polarizers make this difficult, so it is better to estimate it without the polarizers in place. In the Discussion section, we outline a practical approach to adjust the dark hole command to account for the wavefront error introduced by the polarizers. For simplicity, we consider only estimating the real and imaginary parts of the cross field, $\boldsymbol{f}_{y}(\boldsymbol{0})$. When we apply the linear polarizers as described in Sec. 3.1, the measured intensity in the detector pixels with probe command $\boldsymbol{c}_{i}$ is:
@@ 6 Simulations and Results
All optical propagation computations were carried out using the VirtualLab Fusion (VLF) software from LightTrans, Inc. The simulation results in this article correspond to a square Lyot-type stellar coronagraph shown schematically in Fig. 7 and with key parameters summarized in Table 1. This coronagraph consists of 3 OAPs. The initial propagation direction of the beam before it encounters OAP1 is in the $+z$ direction. The simulation is configured so that the principal axis of the beam remains in the $x-z$ plane after each OAP encounter. With this arrangement, the $y-$axis is invariant throughout the system. Assuming no polarizers in the system, the consequence of this invariance is that $y-$polarized light in the input beam results in $y-$polarized light in the detector plane, weak polarization effects due to the finite $f$/# not withstanding. Similarly, if the light in the input beam is $x-$polarized it must also be $x-$polarized in the detector plane since it must be orthogonal to the global $y-$ and local $z-$directions (again, weak polarization effects due to the finite $f$/# not withstanding). Thus, the $A$ and $B$ directions discussed in Sec. 3 correspond directly to the $x$ and $y$ axes in the input beam. To construct the $\tilde{\boldsymbol{D}_{x}}$ and $\tilde{\boldsymbol{D}_{y}}$ matrices introduced in Sec. 4, VLF carried out 1089 propagations-one for each DM actuator-through the optical system. For each of these propagations, the input beam was $x-$polarized and set to one of the DM influence functions with an amplitude of unity. In the detector plane, the corresponding dominant field values, corresponding to a column of $\tilde{\boldsymbol{D}_{x}}$, were the “$E_{x}$” values of the VLF output at each pixel, and the cross field values, corresponding to a column of $\tilde{\boldsymbol{D}_{y}}$, were the “$E_{y}$” values.
@@ 7 Conclusions and Discussion
Disclosures
The author has no conflicts of interest.
Data Availability Statement
The Python codes and supporting data to reproduce the results given in this article are publicly available on GitHub at https://github.com/ComputationalAstrologer/Optics/tree/master/EFC